OpenStudy (bossimbacon):

The altitude of a plane as it begins to descend is shown by the linear equation a = 28,500 – 1800m, where a is the plane’s altitude, in feet, and m is minutes. What is the plane’s altitude when m = 9? @M0j0jojo

OpenStudy (anonymous):

Plug in 9 for m a=28500-1800(9) a=28500-16200 a=12300 Feet